What is the volume in litres of 6 × 1023 gaseous molecules at standard temperature and pressure?


1
Expert's answer
2021-03-05T03:32:18-0500

Solution:

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Moles of gas = (6×1023 gaseous molecules) × (1 mol / 6.022×1023 molecules) = 0.996 mol of gas


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 0.996 mol of gas occupies:

(0.996 mol × 22.4 L) / 1 mol = 22.32 L of gas

0.996 mol of gas occupies a volume equal to 22.32 L.


Answer: The volume of 6×1023 gaseous molecules is 22.32 liters.